Flobberworms have two possibilities for their skin: smooth and scaly. Smooth skin is dominant to scaly. Hagrid mates two smooth-skinned flobberworms, both of which had with one scaly-skinned parent. What proportion of their offspring will have smooth skin?
Answer #1

Smooth skin = SS, Ss

Scaly skin = ss

Genotype of hagrid mates = Ss (not SS, because of them had a parent with scaly skin, ss)

Ss × Ss

Smooth skin = 3/4

Scaly skin = 1/4

Gametes S s
S SS smooth Ss smooth
s Ss smooth ss scaly

Answer is 3/4

Please rate high.